# DRExplainer: Quantifiable Interpretability in Drug Response Prediction with Directed Graph Convolutional Network

### v1

@@ 3.5 Case study
Therefore, the two case studies demonstrate that DRExplainer could help find out the novel drug response and reveal its inner mechanism.
@@ 5 CRediT authorship contribution statement
Haoyuan Shi: Writing - original draft, Writing - review & editing, Methodology, Formal analysis, Visualization. Tao Xu: Investigation, Data curation, Validation, Visualization. Xiaodi Li: Data curation, Validation, Visualization. Qian Gao: Data curation, Validation, Visualization. Junfeng Xia: Supervision. Zhenyu Yue: Supervision, Project administration, Writing - review & editing.
@@ 7 Acknowledgements
This work was supported by the National Natural Science Foundation of China (grant number 62102004) and the Introduction and Stabilization of Talent Project of Anhui Agricultural University (grant number yj2019-32).

### v2

@@ 3.5 Case study
Therefore, the two case studies demonstrate that DRExplainer could help find out the novel drug response and reveal its underlying mechanism. By aligning predictions with existing biological knowledge, the model highlighted its potential to guide personalized treatment strategies. Moreover, DRExplainer offered critical insights that can assist clinical decision-making and avoid ineffective therapies. These examples illustrate how the predictions of the model can directly influence real-world clinical practices and improve patient outcomes.
@@ 4 Conclusion
The implications of this study extend from theoretical advancements in drug response prediction to practical benefits for clinical and preclinical applications. First, by providing interpretable predictions, DRExplainer offers reliable insights to support clinical decision-making and presents a novel approach to advancing the clinical adoption of deep learning models in oncology. Second, the ability of DRExplainer to predict both sensitivity and resistance response can streamline the preclinical evaluation of potential therapies, reducing the time and cost associated with developing new drugs. Lastly, this work advances precision medicine by mitigating the trial-and-error approach in cancer treatment, enabling more patients to receive effective, personalized therapies tailored to their unique profiles.
@@ 5 CRediT authorship contribution statement
Haoyuan Shi: Writing - original draft, Writing - review & editing, Methodology, Formal analysis, Visualization. Tao Xu: Investigation, Data curation, Validation, Visualization. Xiaodi Li: Data curation, Validation, Visualization. Qian Gao: Data curation, Validation, Visualization. Zhiwei Xiong: Formal analysis, Writing - review & editing. Junfeng Xia: Supervision. Zhenyu Yue: Supervision, Project administration, Writing - review & editing.
@@ 7 Acknowledgements
This work was supported by the grants from the National Natural Science Foundation of China (62472005, 62102004), and the Anhui Province Excellent Young Teacher Training Project (YQYB2024007).